# iWISDM: Assessing instruction following in multimodal models at scale

### v1

@@ Abstract
The ability to perform complex tasks from detailed instructions is a key to many remarkable achievements of our species. As humans, we are not only capable of performing a wide variety of tasks but also very complex ones that may entail hundreds or thousands of steps to complete. Large language models and their more recent multimodal counterparts that integrate textual and visual inputs have achieved unprecedented success in performing complex tasks. Yet, most existing benchmarks are largely confined to single-modality inputs (either text or vision), narrowing the scope of multimodal assessments, particularly for instruction-following in multimodal contexts. To bridge this gap, we introduce the instructed-Virtual VISual Decision Making (iWISDM) environment engineered to generate a limitless array of vision-language tasks of varying complexity. Using iWISDM, we compiled three distinct benchmarks of instruction following visual tasks across varying complexity levels and evaluated several newly developed multimodal models on these benchmarks. Our findings establish iWISDM as a robust benchmark for assessing the instructional adherence of both existing and emergent multimodal models and highlight a large gap between these models’ ability to precisely follow instructions with that of humans.
@@ 1 Introduction
A typical day in most people’s lives may involve hundreds or thousands of tasks, most of which are performed without explicit attention. Just in between getting up and getting to work, one may have already performed 5-15 tasks (taking a shower, shaving, making coffee, getting dressed, etc.). Teaching artificial agents to perform similar seemingly mundane tasks have proven to be an extremely difficult computational problem (Konar, 2018). The challenge becomes more apparent when one realizes that each of these seemingly mundane tasks such as making coffee involve tens of steps/actions (Figure 1). The challenge becomes increasingly more significant once we consider more complex tasks such as operating a device or assembling a piece of furniture from its instruction manual. And yet, these tasks are performed proficiently by most individuals in most situations.
@@
Figure 1: Illustration of coffee making task as an example real world compositionally constructed task. a) Cartoon depiction of a kitchen with typical objects within. The dialogue boxes highlight the relevant properties of the grinder, coffee bag, and boiler; b) Pseudocode detailing the coffee-making task, encompassing three subtasks – obtaining coffee powder, brewing the coffee, and returning the coffee; c) Graphical representation of the three subtasks as computational graphs. Blue text highlights the properties associated with each operator; d) An abstraction of the real-life coffee-making task as a task within the iWISDM environment. Frame sequences are generated using the task graphs illustrated in c; e) Graphic depiction of object properties actively in use (top) and subtasks (bottom). The colored bars in the upper panel signify the persistence of each object property during the execution of coffee-making task. The lower panel depicts the interrelations between the subtasks and various temporal compositional operations.
@@
However, existing benchmarks for assessing such models face several shortcomings: (1) Most multimodal benchmarks like FLEURS (audio based, (Conneau et al., 2023)) and VATEX (video based, (Wang et al., 2019)) are still unimodal in their inputs, and do not permit detailed assessment of models’ capacity to integrate information across modalities towards task goals. (2) Visual Question-Answering (VQA) datasets like VQAv2 (Goyal et al., 2017) and CLEVR (Johnson et al., 2017) assess reasoning with visual information in static images without addressing temporal information integration and sequential decision-making. (3) Open-ended learning environments such as XLand (Team et al., 2021), Crafter (Hafner, 2021), and Minecraft (Guss et al., 2019) have been utilized for training reinforcement learning agents, but it remains unclear whether and how they can be adapted to benchmark LMMs. (4) To the best of our knowledge, none of the existing benchmarks specifically assess models’ ability to precisely follow instructions in the context of decision making tasks, an important measure of reliability and trustworthiness. Despite its importance, conducting such assessments have been particularly challenging because of the difficulty of collecting samples of multi-step tasks with ground truth information. (5) More recent benchmarks such as MME (Fu et al., 2023), Mmbench (Liu et al., 2023c) and MMvet (Yu et al., 2023) cover a wide range of cognitive tasks that start to adopt manually-generated or GPT-powered responses. However, those benchmarks are difficult to scale, which makes them inconvenient when investigating scaling properties of LMMs. In addition, benchmarks such as OwlEval (Ye et al., 2023) and LVLMeHub (Xu et al., 2023) rely on subjective human responses that often show high variability across individuals.
@@
To close this gap, we designed instructed-Virtual Visual Decision Making (iWISDM), a virtual environment that enables procedural generation of complex, multi-step decision making tasks that test an agent’s capacity to process visual information guided by natural language instructions. iWISDM builds on the compositional nature of natural behavior (Barker, 1963) and the fact that complex tasks are often compositionally constructed by combining smaller task units together in time. We thus developed a framework which allows instantiating visual decision making tasks as computational graphs that could be combined together in different ways to construct a virtually infinite number of tasks with varying complexity. The code of iWISDM is available on GitHub at https://github.com/BashivanLab/iWISDM.git.
@@
    We introduce iWISDM, a virtual environment for procedural generation of limitless visual decision making tasks accompanied by natural language instructions.
@@
    We test several recently developed LMMs as well as human subjects on these benchmarks and identify a notable shared weakness across existing LMMs compared to humans in their ability to precisely follow user instructions in the context of visual decision making tasks.
@@ 2.1 Large Multimodal Models
A range of methods and models have been developed to enhance the capabilities of LMMs. Early approaches uses a frozen object detector for visual feature extraction (Chen et al., 2020; Li et al., 2020; Zhang et al., 2021) while LiT (Zhai et al., 2022) borrowed a frozen pretrained image encoder from CLIP (Radford et al., 2021). More recently, Frozen (Tsimpoukelli et al., 2021) and Flamingo (Alayrac et al., 2022) have adopted an image-to-text generation approach, prompting the language model to generate text based on an input image. However, in BLIP-2 (Li et al., 2023b), it was shown that this approach is not adequate for overcoming the modality gap and instead proposed a Querying Transformer (Q-Former), acting as a visual resampler, and a two-stage bootstrapping pretraining method, which led to models that outperformed Flamingo80B (Alayrac et al., 2022) in zero-shot VQAv2 tasks with fewer trainable parameters. The Q-former architecture was also adopted by later work such as InstructBLIP (Dai et al., 2023) and Qwen-VL (Bai et al., 2023).
@@ 3 Methodology
We developed iWISDM to facilitate the generation of a diverse range of sequential visual reasoning tasks, which vary in complexity and require minimal user intervention. iWISDM encompasses a broad spectrum of tasks that engage executive functions such as inhibition of action, working memory, attentional set, task switching, and schema generalization. These functions are traditionally associated with the prefrontal cortex, a critical area for advanced cognitive processes in the brain (Fuster, 2015; Sun et al., 2023). Notably, the iWISDM task space is deliberately designed to accommodate a multitude of classic working memory and decision-making tasks commonly employed in neuroscience and cognitive science research (Rigotti et al., 2013; Goldman-Rakic, 1992; Fuster, 2009). In the following sections, we begin by detailing the key functionalities of the iWISDM environment, followed by an in-depth description of its design and the implementation of its constituent components.
@@ 3.1.1 Task graph construction
  • Functional Operators: Select: This operator defines stimuli based on three criteria: when, where, and what. “When” refers to the specific frame from which the stimuli originate. “Where”indicates the location of the stimuli within the frame. “What” depends on the particular dataset from which the stimuli are derived. For example, in our ShapeNet environment, the stimuli have three attributes: category (such as car versus plane), identity (which specific car), and view angle (the angle from which the stimulus is rendered). Select operators that have no downstream connections are the terminal nodes of the graph. Conversely, its potential downstream operators may include any Get* functional operator. Switch: Based on the output action of a boolean task, this operator connects the logic to one of two possible paths. Its compulsory downstream connection must be a boolean operator, while its typical upstream connections are subtasks graphs. Get*: This group of operators is responsible for fetching specific properties of stimuli, such as category, location, or identity, exemplified by operators like GetCategory, GetLocation, and GetIdentity. Its direct downstream connection is always a Select operator, and its upstream connections can be any boolean operator. CONST: The simplest form of operator, CONST represents a fixed value. It is often used as a downstream connection for boolean operators that compare attributes.
  • Boolean Operators: Exist: Paired with a specific property value (e.g., ’Desk’), this operator poses the question of whether an object with the property (’Desk’) exists. It generates a boolean output and can function as an action generator or as a downstream connection for Switch and other boolean operators. And, IsSame, NotSame, Or: These boolean operators process inputs from two boolean operators to produce a boolean outcome. They are critical for constructing logical conditions within the task graph.
Using these operators, iWISDM provides two possible modes of constructing task graphs: user-specified and automatic.
@@
In addition to above hyperparameters, in AutoTask mode, the allowed task structure is further constrained by the permitted connectivity for various operators (e.g. And operators must be followed by other boolean operators such as IsSame, NotSame, And, Or). A default operator connectivity is defined for all existing operators in our core build, but new connectivity rules could easily be added for any new user-defined operators. This is done through a Python dictionary, which details the allowed input and output operators for additional operators. By specifying these hyperparameters, iWISDM autonomously generates random runnable task graphs derived from the predefined task space. Each resultant task graph can then be used to generate specific trials.
To assure each task graph would comply with the connectivity rules between operators, we follow a backwards initialization process during AutoTask (Figure A1). The task generation process starts from the root node and descends recursively. For each current node/operator $n$ in the graph, its downstream operators $C_{n}$ are randomly sampled based on the connectivity rules. As the graph depth approaches the specified maximum depth, the permissible operators with the shortest possible subtask depth are sampled into $C_{n}$. For instance in our core build, if $n$ is the And operator, then only IsSame, NotSame are sampled since they have shorter subgraph depths than And, Or. Through this procedure, iWISDM AutoTask facilitates sampling from diverse task spaces with varying degrees of complexity specified by the user. The utilization of the connectivity rule dictionary and Switch operator guarantees the logical and feasible nature of the generated tasks, providing researchers with an extensive pool of tasks for investigation and exploration.
@@ 3.1.2 Node Initialization
Figure 2: Illustration of various temporal compositional structures used in iWISDM and their application in instantiating classic decision making tasks. a) Four distinct temporal composition operations including Queue: where one task is completed before starting the next; Overlap: where two or more tasks share common information; Interleave: depicting the interwoven acquisition of information related to different tasks and; Condition: where the execution of a subsequent task depends on the outcome of the preceding one. An example compositional task consisting of multiple temporal relationships is shown. Frames are colored differently to highlight distinct tasks, with multicolored frames indicating shared information across tasks; b) Three classic cognitive tasks are exemplified: Delayed Match to Sample; 2-Back; and Contextual Decision Making.
@@ 3.1.3 Task trial instantiation
  • Frame sequence. The frame sequence consists of a array of images. They display the visual information at each time step of the trial. Each frame is accompanied by a dictionary that contains the object properties within that frame. Images are stored as PNG files.
@@
Figure 3: Description of operators and task graph examples in the core build of iWISDM. a) Two main types of operators are considered. Functional Operators: Select operators retrieve stimuli according to specified attributes such as time and location; Switch operators accept boolean values and direct the task logic to one of two different subtasks; GetAttribute operators take stimuli as input and output the corresponding attribute values of those stimuli. Boolean Operators comprise Exit, And, Or, IsSame, and NotSame. These operators take boolean values as inputs and produce another boolean value based on their boolean logic. b-d) Using the operators defined in panel a, we can create tasks following pre-specified rules. Panels b, c, and d display the task graphs for the three typical cognitive tasks illustrated in Figure 2b, demonstrating how these operators are interconnected to define the structures of these specific tasks. For instance, the task instruction for panel b and c are: “category of object 1 equals category of object 2?”, “if category of object 1 equals category of object 3, then category of object 2 equals category of object 3, else category of object 2 equals category of object 4?’. The task instruction for panel c is similar to panel b, but repeated across time.
@@ 3.2 Main Features
At its core, iWISDM is designed to be scalable and extensible. To do so, we adopted a modular framework in which task rules are constructed compositionally by combining functional and boolean operators. The combination of these operators give rise to distinct tasks. Likewise, task spaces (i.e. collections of instantiable task graphs) are spanned by specifying the set of allowable operators and operator-operator connection rules (see section 3.1 for detailed definitions). We detail iWISDM’s main features below.
Compositionality. Real world tasks are fundamentally compositional, as most tasks can be readily decomposed into sets of simpler subtasks that involve fewer sensory observations and cognitive operations. There are two crucial facets of compositionality that need to be considered: logical and temporal, both of which are common in daily human behavior, and allow individuals to efficiently handle complex tasks and adapt to dynamic environments. Cognitive processes involving task decomposition (i.e. breaking tasks into subtasks) and temporal combination of decision rules (i.e. combining the outcomes of subtasks) are fundamental to our ability to navigate the world around us.
  • Logical: An agent’s action can be viewed as a function of sensory observations, internalized world knowledge, prior actions, and its objectives (Ha & Schmidhuber, 2018). Yet, as outlined in the introduction, the decision-making process frequently decomposes into sub-decisions and information processing steps that are temporally constrained and required fewer observations. We define logical compositionality as how decisions can be combined hierarchically through boolean operators (i.e. And, Or, etc) and functional operators (i.e. Switch operator that asks for if…then…). As an example, consider the contextual decision making task (ctxDM, see Figure 2b or Figure 3d). In the task in Figure 3d, the subject needs to first compare the category of the objects in the first and third frames. If their categories are the same, then compare the category of the objects from the second and third frames. Otherwise, compare the category of the objects from the second and fourth frames. For this task, the task rule of the second task is conditioned upon the result of the first task, which makes a logical composition. As the depth and the total number of operators involved in the task grows, we can compose more complex logical structures.
@@
Vast Task Space. Another important feature of iWISDM is the vastness of its task space. This featurenaturally extends from the compositionality that is inherent to iWISDM’s task generation procedure. The ability to produce a large number of distinct tasks is critically important for the robust evaluation of large multimodal models that are trained on increasing volume of information from the web. Moreover, the vast number of instantiable tasks in iWISDM provides an opportunity for training or fine-tuning large multimodal models to improve their ability to follow instructions in a vision-language context.
Natural Language Task Instruction. Natural language provides a rich and convenient way of communicating complex information to biological or artificial agents. It has been shown that improvements on language understanding in LLMs directly enhances their generality (performing many tasks) and adaptation (0-shot generalization)Brown et al. (2020); Radford et al. (2019). Due to their capacity to compress enormous knowledge bases, these models have been useful in various applications where traditionally human supervision had been necessary Shah et al. (2023); Dasgupta et al. (2023). Perhaps for similar reasons, natural language input constitutes the core of most existing multimodal models and they are heavily trained on large text corpuses among others. For this reason, in iWISDM, each task is accompanied by a simplified natural language instruction (see examples A1). When completing complex tasks, the instruction communicates first the task structure in terms of upcoming observations, then the task rules that determine the relationship between observations and actions.
Automatic Task Generation. In contrast to prior virtual environments tailored for the purposes of cognitive or neuroscience investigations (Molano-Mazon et al., 2022), iWISDM can not only generate hand-crafted cognitive tasks such as classical decision making tasks (e.g. contextual decision making and n-back, 3), but also allow procedural task generation from a pre-specified task space defined by a small set of hyperparameters (See Section 3.2 for details).
Customizability and Extensibility. We envision the future of iWISDM as a framework that will be continuously developed and expanded by the larger community of machine learning scientists, cognitive scientists, and neuroscientists. For this reason, we have designed iWISDM to be highly customizable and extendable in several ways.
  • Visual inputs. iWISDM also allows any natural image set to be used as the stimulus set. In our core build, we use 2D projection of 3D object models from the ShapeNet dataset (Chang et al., 2015), where each stimuli is defined by a parameter vector consisting of category, identity, pose angle, location. We provide a template that allows users to seamlessly import alternative stimulus datasets.

### v3

@@ Abstract
The ability to perform complex tasks from detailed instructions is a key to many remarkable achievements of our species. As humans, we are not only capable of performing a wide variety of tasks but also very complex ones that may entail hundreds or thousands of steps to complete. Large language models and their more recent multimodal counterparts that integrate textual and visual inputs have achieved unprecedented success in performing complex tasks. Yet, most existing benchmarks are largely confined to single-modality inputs (either text or vision), narrowing the scope of multimodal assessments, particularly for instruction-following in multimodal contexts. To bridge this gap, we introduce the instructed-Virtual VISual Decision Making (iWISDM) environment engineered to generate a limitless array of vision-language tasks of varying complexity. Using iWISDM, we compiled three distinct benchmarks of instruction following visual tasks across varying complexity levels and evaluated several newly developed multimodal models on these benchmarks. Our findings establish iWISDM as a robust benchmark for assessing the instructional adherence of both existing and emergent multimodal models and highlight a large gap between these models’ ability to precisely follow instructions with that of humans. The code of iWISDM is available on GitHub at https://github.com/BashivanLab/iWISDM.
@@ 1 Introduction
A typical day in most people’s lives involves hundreds or thousands of tasks. Most of which are performed without explicit attention. Just in between getting up and getting to work, one may have already performed 5-15 tasks (taking a shower, shaving, making coffee, getting dressed, etc.). Teaching artificial agents to perform similar seemingly mundane tasks has proven to be an extremely difficult computational problem (Konar, 2018). The challenge becomes more apparent when one realizes that each of these seemingly mundane tasks such as making coffee involve tens of steps/actions (Figure 1). The challenge becomes increasingly more significant once we consider more complex tasks such as operating a device or assembling a piece of furniture from its instruction manual. And yet, these tasks are performed proficiently by most individuals in most situations.
@@
Figure 1: Illustration of coffee making task as an example real world compositionally constructed task. a) Cartoon depiction of a kitchen with typical objects within. The dialogue boxes highlight the relevant properties of the grinder, coffee bag, and boiler; b) Pseudocode detailing the coffee-making task, encompassing three subtasks – obtaining coffee powder, brewing the coffee, and returning the coffee; c) Graphical representation of the three subtasks as computational graphs. Blue text highlights the properties associated with each operator; d) An abstraction of the real-life coffee-making task as a task in iWISDM. Frame sequences are generated using the task graphs illustrated in c; e) Graphic depiction of object properties actively in use (top) and subtasks (bottom). Coloured bars in the upper panel signify the persistence of each object property during the execution of coffee-making task. Lower panel depicts the interrelations between subtasks and various temporal compositional operations.
@@
However, existing benchmarks for assessing such models face several shortcomings: (1) Most multimodal benchmarks like FLEURS (audio-based, (Conneau et al., 2023)) and VATEX (video-based, (Wang et al., 2019)) are still unimodal in their inputs, and do not permit detailed assessment of models’ capacity to integrate information across modalities towards task goals. (2) Visual Question-Answering (VQA) datasets like VQAv2 (Goyal et al., 2017) and CLEVR (Johnson et al., 2017) assess reasoning with visual information in static images without addressing temporal information integration and sequential decision-making. (3) Open-ended learning environments such as XLand (Team et al., 2021), Crafter (Hafner, 2021), and Minecraft (Guss et al., 2019) have been utilized for training reinforcement learning agents, but it remains unclear whether and how they can be adapted to benchmark LMMs. (4) To the best of our knowledge, none of the existing benchmarks specifically assess models’ ability to precisely follow instructions in the context of decision making tasks, an important measure of reliability and trustworthiness. Despite its importance, conducting such assessments have been particularly challenging because of the difficulty of collecting samples of multi-step tasks with ground truth information. (5) More recent benchmarks such as MME (Fu et al., 2023), Mmbench (Liu et al., 2023c) and MMvet (Yu et al., 2023) cover a wide range of cognitive tasks that start to adopt manually-generated or GPT-powered responses. However, those benchmarks are difficult to scale, which makes them inconvenient when investigating the scaling properties of LMMs. In addition, benchmarks such as OwlEval (Ye et al., 2023) and LVLMeHub (Xu et al., 2023) rely on subjective human responses that often show high variability across individuals.
@@
To close this gap, we designed instructed-Virtual Visual Decision Making (iWISDM), a virtual environment that enables procedural generation of complex, multi-step decision making tasks that test an agent’s capacity to process visual information guided by natural language instructions. iWISDM builds on the compositional nature of natural behaviour (Barker, 1963) and the fact that complex tasks are often compositionally constructed by combining smaller task units in time. We thus developed a framework which allows instantiating visual decision-making tasks as computational graphs that could be combined logically and temporally to construct a virtually infinite number of tasks with varying complexity. The code of iWISDM is available on GitHub at https://github.com/BashivanLab/iWISDM.
@@
    We introduce iWISDM, a virtual environment for procedural generation of limitless visual decision-making tasks accompanied by natural language instructions.
@@
    We test several recently developed LMMs as well as human subjects on these benchmarks and identify a notable shared weakness across existing LMMs compared to humans in their ability to precisely follow user instructions in the context of visual decision-making tasks.
@@ 2.1 Large Multimodal Models
A range of methods and models have been developed to enhance the capabilities of LMMs. Early approaches use a frozen object detector for visual feature extraction (Chen et al., 2020; Li et al., 2020; Zhang et al., 2021) while LiT (Zhai et al., 2022) borrowed a frozen pretrained image encoder from CLIP (Radford et al., 2021). More recently, Frozen (Tsimpoukelli et al., 2021) and Flamingo (Alayrac et al., 2022) have adopted an image-to-text generation approach, prompting the language model to generate text based on an input image. However, in BLIP-2 (Li et al., 2023b), it was shown that this approach is not adequate for overcoming the modality gap and instead proposed a Querying Transformer (Q-Former), acting as a visual resampler, and a two-stage bootstrapping pretraining method, which led to models that outperformed Flamingo80B (Alayrac et al., 2022) in zero-shot VQAv2 tasks with fewer trainable parameters. The Q-former architecture was also adopted by later works such as InstructBLIP (Dai et al., 2023) and Qwen-VL (Bai et al., 2023).
@@ 3 Methodology
We developed iWISDM to facilitate the generation of a diverse range of sequential visual reasoning tasks, which vary in complexity and require minimal user intervention. iWISDM encompasses a broad spectrum of tasks that engage executive functions such as inhibition of action, working memory, attentional set, task switching, and schema generalization. These functions are traditionally associated with the prefrontal cortex, a critical area for advanced cognitive processes in the brain (Fuster, 2015; Sun et al., 2023). Notably, the iWISDM task space is designed to accommodate classic working memory and decision-making tasks commonly employed in neuroscience and cognitive science research (Rigotti et al., 2013; Goldman-Rakic, 1992; Fuster, 2009). In the following sections, we begin by detailing the key functionalities of the iWISDM environment, followed by an in-depth description of its design and the implementation of its constituent components.
@@ 3.1.1 Task graph construction
  • Functional Operators: Select: This operator defines stimuli based on three criteria: when, where, and what. “When” refers to the specific frame from which the stimuli originate. “Where” indicates the location of the stimuli within the frame. “What” depends on the particular dataset from which the stimuli are derived. For example, in our ShapeNet environment, the stimuli have three attributes: category (such as car versus plane), identity (which specific car), and view angle (the angle from which the stimulus is rendered). Select operators that have no downstream connections are the terminal nodes of the graph. Conversely, its potential downstream operators may include any Get* functional operator. Switch: Based on the output action of a boolean task, this operator connects the logic to one of two possible paths. Its compulsory downstream connection must be a boolean operator, while its typical upstream connections are subtasks graphs. Get*: This group of operators is responsible for fetching specific properties of stimuli, such as category, location, or identity, exemplified by operators like GetCategory, GetLocation, and GetIdentity. Its direct downstream connection is always a Select operator and its upstream connections can be any boolean operator. CONST: The simplest operator, CONST represents a fixed value. It is often used as a downstream connection for boolean operators that compare attributes.
  • Boolean Operators: Exist: Paired with a specific property value (e.g., ’Desk’), this operator poses the question of whether an object with the property (’Desk’) exists. It generates a boolean output and functions as an action generator or downstream connection for boolean operators. And, IsSame, NotSame, Or: These boolean operators process inputs from two boolean operators to produce a boolean outcome. They are critical in constructing logical conditions within the task graph.
With these operators, iWISDM provides two modes for constructing task graphs: user-specified and automatic.
@@
In addition to the above hyperparameters, in AutoTask mode, the allowed task structure is further constrained by the permitted connectivity for various operators (e.g. And operators must be followed by other boolean operators such as IsSame, NotSame, And, Or). A default operator connectivity is defined for all existing operators in our core build, but new connectivity rules could easily be added for any new user-defined operators. This is done through a Python dictionary, which details the allowed input and output operators for additional operators. By specifying these hyperparameters, iWISDM autonomously generates random runnable task graphs derived from the predefined task space. Each resultant task graph can then be used to generate specific trials.
To ensure each task graph would comply with the connectivity rules between operators, we follow a backwards initialization process during AutoTask (Figure A1). The task generation process starts from the root node and descends recursively. For each current node/operator $n$ in the graph, its downstream operators $C_{n}$ are randomly sampled based on the connectivity rules. As the graph depth approaches the specified maximum depth, the permissible operators with the shortest possible subtask depth are sampled into $C_{n}$. For instance in our core build, if $n$ is the And operator, then only IsSame, NotSame are sampled since they have shorter subgraph depths than And, Or. Through this procedure, iWISDM AutoTask facilitates sampling from diverse task spaces with varying degrees of complexity specified by the user. The utilization of the connectivity rule dictionary and Switch operator guarantees the logical and feasible nature of the generated tasks, providing researchers with an extensive pool of tasks for investigation and exploration.
@@ 3.1.2 Node Initialization
Figure 2: Illustration of various temporal compositional structures used in iWISDM and their application in instantiating classic decision-making tasks. a) Four distinct temporal composition operations including Queue: where one task is completed before starting the next; Overlap: where two or more tasks share common information; Interleave: depicting the interwoven acquisition of information related to different tasks and; Condition: where the execution of a subsequent task depends on the outcome of the preceding one. An example compositional task consisting of multiple temporal relationships is shown. Frames are coloured differently to highlight distinct tasks, with multicoloured frames indicating shared information across tasks; b) Three classic cognitive tasks are exemplified: Delayed Match to Sample; 2-Back; and Contextual Decision Making.
@@ 3.1.3 Task trial instantiation
  • Frame sequence. The frame sequence consists of an array of images. They display the visual information at each time step of the trial. Each frame is accompanied by a dictionary that contains the object properties within that frame. Images are stored as PNG files.
@@
Figure 3: Description of operators and task graph examples in the core build of iWISDM. a) Two main types of operators are considered. Functional Operators: Select operators retrieve stimuli according to specified attributes such as time and location; Switch operators accept boolean values and direct the task logic to one of two different subtasks; GetAttribute operators take stimuli as input and output the corresponding attribute values of those stimuli. Boolean Operators comprise Exit, And, Or, IsSame, and NotSame. These operators take boolean values as inputs and produce another boolean value based on their boolean logic. b-d) Using the operators defined in panel a, we can create tasks following pre-specified rules. Panels b, c, and d display the task graphs for the three typical cognitive tasks illustrated in Figure 2b, demonstrating how these operators are interconnected to define the structures of these specific tasks. For instance, the task instruction for panel b and c are: “category of object 1 equals category of object 2?”, “if category of object 1 equals category of object 3, then category of object 2 equals category of object 3, else category of object 2 equals category of object 4?’. The task instruction for panel c is similar to panel b but repeated across time.
@@ 3.2 Main Features
At its core, iWISDM is designed to be scalable and extensible. To do so, we adopted a modular framework in which task rules are constructed compositionally by combining functional and boolean operators. The combination of these operators gives rise to distinct tasks. Likewise, task spaces (i.e. collections of instantiable task graphs) are spanned by specifying the set of allowable operators and operator-operator connection rules (see section 3.1 for detailed definitions). We detail iWISDM’s main features below.
Compositionality. Real-world tasks are fundamentally compositional, as most tasks can be readily decomposed into sets of simpler subtasks that involve fewer sensory observations and cognitive operations. Two crucial facets of compositionality need to be considered: logical and temporal, both of which are common in daily human behavior and allow individuals to efficiently handle complex tasks and adapt to dynamic environments. Cognitive processes involving task decomposition (i.e. breaking tasks into subtasks) and temporal combination of decision rules (i.e. combining the outcomes of subtasks) are fundamental to our ability to navigate the world around us.
  • Logical: An agent’s action can be viewed as a function of sensory observations, internalized world knowledge, prior actions, and its objectives (Ha & Schmidhuber, 2018). Yet, as outlined in the introduction, the decision-making process frequently decomposes into sub-decisions and information-processing steps that are temporally constrained and require fewer observations. We define logical compositionality as how decisions can be combined hierarchically through boolean operators (i.e. And, Or, etc) and functional operators (i.e. Switch operator that asks for if…then…). As an example, consider the contextual decision-making task (ctxDM, see Figure 2b or Figure 3d). In the task in Figure 3d, the subject needs to first compare the category of the objects in the first and third frames. If their categories are the same, then compare the category of the objects from the second and third frames. Otherwise, compare the category of the objects from the second and fourth frames. For this task, the task rule of the second task is conditioned upon the result of the first task, which makes a logical composition. As the depth and the total number of operators involved in the task grows, we can compose more complex logical structures.
@@
Vast Task Space. Another important feature of iWISDM is the vastness of its task space. This feature naturally extends from the compositionality that is inherent to iWISDM’s task generation procedure. The ability to produce a large number of distinct tasks is critically important for the robust evaluation of large multimodal models that are trained on increasing volumes of information from the web. Moreover, the vast number of instantiable tasks in iWISDM provides an opportunity for training or fine-tuning large multimodal models to improve their ability to follow instructions in a vision-language context.
Natural Language Task Instruction. Natural language provides a rich and convenient way of communicating complex information to biological or artificial agents. It has been shown that improvements on language understanding in LLMs directly enhance their generality (performing many tasks) and adaptation (0-shot generalization)Brown et al. (2020); Radford et al. (2019). Due to their capacity to compress enormous knowledge bases, these models have been useful in various applications where traditionally human supervision had been necessary Shah et al. (2023); Dasgupta et al. (2023). Perhaps for similar reasons, natural language input constitutes the core of most existing multimodal models and they are heavily trained on large text corpora among others. For this reason, in iWISDM, each task is accompanied by a simplified natural language instruction (see examples A1). When completing complex tasks, the instruction communicates first the task structure in terms of upcoming observations, then the task rules that determine the relationship between observations and actions.
Automatic Task Generation. In contrast to prior virtual environments tailored for cognitive or neuroscience investigations (Molano-Mazon et al., 2022), iWISDM can not only generate hand-crafted cognitive tasks such as classical decision-making tasks (e.g. contextual decision-making and n-back, 3), but also allow procedural task generation from a pre-specified task space defined by a small set of hyperparameters (See Section 3.2 for details).
Customizability and Extensibility. We envision the future of iWISDM as a framework that will be continuously developed and expanded by the larger community of machine learning scientists, cognitive scientists, and neuroscientists. For this reason, we have designed iWISDM to be highly customizable and extendable in task operators, visual inputs, and stimulus properties (for detailed discussions, see Appendix A.1.1).